OpenStudy (anonymous):

this proeblem seems so easy but yet so hard :( can someone please help me? if 2x-3(x+4)=-5, then x=??

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (radar):

Remove the parenthesis: 2x-3x-12=-5

OpenStudy (radar):

combine similar terms -x-12=-5 Add 12 to both sides -x-12+12=-5+12 this was done to get x by itself on one side of the equal sign. -x=7 Oh! you want x positive........multiply both sides by -1 -x(-1)=7(-1) x=-7